# Excitable Nonlinear Opinion Dynamics (E-NOD) for Agile Decision-Making

### v2

@@ IV Multi-agent excitable decision-making
where $A\!=\![a_{ik}]$ is the E-NOD network adjacency matrix, capturing both opinion self-reinforcing terms (diagonal elements) and inter-agent opinion exchange terms (out-diagonal elements). We assume homogeneous agents, i.e., all agents have the same $d$, $u_{0}$, $K_{u}$ and $K_{u_{s}}$. Notice that (5) is exactly the networked, distributed version of (4).
The higher dimensional nature of this system makes its dynamics harder to characterize analytically. However, many of its properties can be inferred from our analysis of the one dimensional case, the analysis of the NOD in [1], and/or from numerical simulations. Here we just list a few notable properties of (5) that appear to hold under mild assumptions. A rigorous analysis is left for future works.
The origin $\textbf{0}\!\in\!\mathbb{R}^{n}$ is always a fixed point for the no input ($b_{i}\!=\!0$) case, but its stability depends on the parameter choice. There is a bifurcation of the origin with respect to parameter $u_{0}$ where the origin goes from being stable to unstable at a critical value $u_{0}^{*}$. For appropriate parameters, one or two limit cycles, symmetric with respect to some hyperplane, appear at this bifurcation. These limit cycles have in general longer periods as compared to the single agent case with the same parameters. For values of $u_{0}$ slightly smaller than $u_{0}^{*}$ the system is excitable and all trajectories approach one of two prototypical spiking decision trajectories (i.e., upward or downward spikes), both of which converge back to the origin. At $u_{0}\!=\!u_{0}^{*}$ these two spiking decision trajectories bifurcate into one or two attracting limit cycles.
When the graph induced by $A$ is regular, then the consensus subspace, $\{(u_{s}\!\cdot\!\mathds{1},z\!\cdot\!\mathds{1})\!\mid\!u_{s},z\!\in\!\mathds{R}\}$ with $\mathds{1}\!=\!(1,\ldots,1)\!\in\!\mathds{R}^{N_{a}}$, is forward invariant. The dynamics on this subspace can be reduced to those of the single agent model. If this subspace is stable, the dynamics lead to opinion synchronization.
V Application to agile, social robot navigation
We use the E-NOD (4) to design an agile controller for a social robot navigating around oncoming human movers. Simulations of the resulting behavior can be seen in Fig. 1 and 5 and are animated here.
Consider a robot moving toward a goal at constant speed in a space that may be occupied by other robots and humans. Each object in the 2D space is described by its position and direction of motion. Each robot seeks to avoid collisions with humans and reach its goal. For each robot, we introduce an E-NOD state $z$ ruled by (4) that determines the robot preference for turning left ($z\!>\!0$), right ($z\!<\!0$), or going straight ($z\!=\!0$), relative to its nominal (goal-tracking) control, in order to avoid other movers and obstacles. That is, we embody the E-NOD (4) in the physical robot steering control. For multi-robot coordination, we use the multi-agent E-NOD (5).
For human-robot interaction and collision avoidance, we define a proxy of the human’s (unknown) heading opinion at time $t$, $\hat{z}_{h}(t)\!=\!\tan(\eta_{h}(t))$, where $\eta_{h}(t)$ is a reconstruction at time $t$ of the human heading direction from the robot viewpoint, i.e., the angle between the (measured) human’s heading direction and the direction of the robot-human vector. Each robot steering opinion $z_{i}$ evolves over time $t$ according to (5), where the opinion interaction weight $a_{ik}$ depends on the distance between robots $i$ and $k$. Human heading direction proxy is added as an input $b\!=\!u\!\cdot\!\hat{z}_{h}$. The basal attention of each robot $u_{0}$ is modulated by the proximity and bearing of a robot to a human. As a robot and human move closer to an impending collision, $u_{0}$ increases.
Fig. 1 compares the trajectory and opinion of one robot navigating around one human mover when controlled by NOD (adapted from [11]) or E-NOD (4). Initially the human appears to pass on the robot’s left, so both robots bear right (with $z\!<\!0$). When the human abruptly turns to pass on the robot’s right, NOD lags to adapt and bear left ($z\!>\!0$), whereas E-NOD allows for a quick and agile change in heading direction. Only the E-NOD controlled robot is able to turn left soon enough to navigate around the human mover.
Fig. 5 showcases a set of three robots coupled through multi-agent E-NOD (5) in a leader-follower configuration as they navigate toward a common target and around two approaching humans. The only difference between the simulation in Fig. 5a and Fig. 5b is the sign of the communication network edge weights. In Fig. 5a, observers follow leader opinion. When the leader robot turns to the left to dodge the approaching humans, all robots follow, gracefully reaching the goal. In Fig. 5b, observers do not follow leader opinion. When the leader robot turns left with $z\!>\!0$, the second robot disagrees and forms $z\!<\!0$ to turn right. In further disagreement, the last robot turns left with $z\!>\!0$. The second robot later interacts with another human in its path, and its opinion adapts to $z\!>\!0$, leading to a timely left turn to avoid a collision. These simulatons show the agility of E-NOD and its sequential decision-making power.
Figure 5: Trajectories of social robot leader-follower teams navigating around approaching human movers. (a) Observers follow leader opinion. (b) Observers do not follow leader opinion. Plot of $z$ and $u_{s}$ over time $t$ for each robot in this case.
VI Final Remarks
@@
We presented Excitable Nonlinear Opinion Dynamics (E-NOD) for a single agent and two options and showcased its ability to swiftly form opinions and regulate back to ultrasensitivity. Parameter $K_{u}$ was identified as a cause for NOD becoming too robust, but the self-regulation of E-NOD recovers flexibility. The existence of limit cycles for certain parameter regimes in E-NOD was found and examined. We presented a multi-agent E-NOD model and highlighted its limit cycles and potential for agent synchronization. We illustrated the agility of E-NOD in an application to social robot navigation. We aim to provide analytical guarantees on the onset of periodic spiking in limit cycles. We will also implement E-NOD control on physical robots in social navigation experiments.

### v3

@@ IV-A S-NOD for Multiple Agents and Two Options
where $A\!=\![a_{ik}]\in\mathbb{R}^{N_{a}\times N_{a}}$ is the S-NOD network adjacency matrix, capturing both the strength ($a_{ii}\geq 0$) of a self-reinforcing term and the strength ($|a_{ik}|$) of the influence of the opinion of agent $k$ on the opinion of agent $i$. If $a_{ik}$ is positive (negative), then an opinion of agent $k$ in favor of one of the options influences an opinion of agent $i$ in favor of the same (other) option. We assume homogeneous agents, i.e., all agents have the same $d$, $u_{0}$, $K_{u}$ and $K_{u_{s}}$. S-NOD as presented in (5) is the networked, distributed version of (4).
In Fig. 5 we simulate the opinion dynamics of (5) for three agents in a network when agent 1 receives a constant input $b_{1}>0$ for $t\in[10,40]$. For the loop-free networks of Fig. 5, we see that when the weights $a_{ik}$ are positive (negative) the spiking of agent 1 for option 1 triggers synchronized (anti-synchronized) spiking of agents 2 and 3. Future work will characterize different possible behaviors, e.g., opinion spike (anti)synchronization, for classes of network structures.
IV-B Application to Social Robot Navigation
Figure 6: Trajectories of social robot teams navigating around approaching human movers. Communication network and parameters are as in Fig. 5 with (a) $a_{ik}=+0.1$ and (b) $a_{ik}=-0.1$. (c) Plots of $z$ over time $t$ for the robots. Table I lists performance metrics of the robots.
We use S-NOD (5) to design a decentralized, agile controller for social robots navigating around oncoming human movers in 2D. Each robot has a nominal control that steers it toward its goal by regulating its heading direction through proportional negative feedback. The S-NOD state $z_{i}$ defines the strength of robot $i$’s preference for turning left ($z_{i}\!>\!0$) or right ($z_{i}\!<\!0$). A term mediated by $z_{i}$ is added as positive feedback to the control on steering behavior. This overcomes negative feedback regulation and promotes fast reactive steering when possible collisions with oncoming human movers are imminent. Simulations of the resulting navigation behavior are in Fig. 1 and 6 and animated at https://spikingNOD.github.io.
To anticipate collisions, robot $i$ can estimate its distance ($\rho_{i}$) to a human, bearing angle ($\eta_{i}$) on the human, and angle ($\eta_{h,i}$) between the human’s heading direction and the robot-human vector. Robots can exchange steering opinions over a communication network as in (5a). The robot’s attention grows above its basal level $u_{0}$ as collision risk grows with decreasing $\rho_{i}$ and $\eta_{i}$. This increases the strength of (i) the positive feedback loop of the steering controller, and (ii) the interactions with other robots to achieve coordinated obstacle avoidance. Thus, each robot’s steering opinion deviates from navigating toward its goal dependent on $\rho_{i}$, $\eta_{i}$, $\eta_{h_{i}}$, and on other robot opinions. We let $\hat{z}_{h}(t)\!=\!\tan(\eta_{h}(t))$ be a proxy for the human’s opinion at time $t$ and add $\hat{z}_{h}$ to the term $\sum_{k=1}^{N_{a}}\!a_{ik}z_{k}$ in (5a). Coordination among robots derives from the sign of $a_{ik}$: when $a_{ik}\!>\!0$ ($a_{ik}\!<\!0$), robot $i$ is influenced to make a similar (opposite) steering choice as robot $k$. We let $|a_{ik}|$ decay with growing distance between robots $i$, $k$.
Fig. 1 compares the trajectory and opinion of a robot using NOD (adapted from [18]) and S-NOD (4) models to navigate around a human mover. The S-NOD robot passes the human with a minimum distance of 0.96m and arrives at its goal in 14.4 seconds. Without the return to the sensitive bifurcation point that S-NOD provides, the NOD robot’s opinion change lags as the human makes a sharp turn and the robot experiences a collision (moving closer than 0.3m to the human) less than 6 seconds into the simulation.
Fig. 6 showcases three robots navigating towards a common goal and around two approaching humans using multi-agent S-NOD (5). The communication networks are those in Fig. 5 and the spiking behavior of the robots in Fig. 6c is similar in many ways to the spiking behavior in Fig. 5a and 5b. However, due to the distance dependent $|a_{ik}|$ of Fig. 6b, robot 3 prefers the same opinion as robot 1, whereas due to the constant $|a_{ik}|$ of Fig. 5b, agent 3 prefers the opinion of agent 2. Notably in Fig. 6b, robot 2 switches from an initial right turn (opposite to robots 1 and 3) to avoid human 1 to a left turn to avoid human 2. S-NOD gracefully navigates robot 2 out of consecutive potential collisions, implementing different turning opinions. This embodies the agility and sequential decision-making feature of S-NOD.
V Final Remarks
@@
We presented and analyzed Spiking Nonlinear Opinion Dynamics (S-NOD) for a single agent and two options. We showcased the ability of S-NOD to swiftly form opinions and regulate back to ultrasensitivity. S-NOD provides first-of-its-kind two-dimensional excitable (spiking) dynamics for agile decision-making over two options. We showed how NOD can become too robust, but the self-regulation of S-NOD recovers flexibility. We analyzed existence of limit cycles for certain parameter regimes in S-NOD. We presented S-NOD for multiple agents that communicate opinions over a network and highlighted potential for agent (anti)synchronization. We illustrated S-NOD’s agility in a social robot navigation application and plan to implement on physical robots. We aim to provide analytical guarantees on the onset of periodic spiking in limit cycles, to analyze synchronization patterns for multiple agents, and to generalize to multiple options.